Eleanor is working her way through school. She works two part-time jobs for a total of 2525 hours a week. Job A pays $6.50$6.50 per hour, and Job B pays $7.30$7.30 per hour. How many hours did she work at each job the week that she made $168.90$168.90.
Answer by Boreal(15235)

A+B=25
6.50A+7.30B=168.90
-6.50A-6.50B=-162.50
0.80B=6.40
B=8 hours@$7.30=$58.40
A=17 hours@6.50=$110.50
They add to $168.90